OpenStudy (anonymous):

A random variable 'X' has the following probability x 0 1 2 3 4 5 6 7 P(X) 0 k 2k 2k 3k k^2 2k^2 7k^2 + k

OpenStudy (anonymous):

We know that sum of all probabilits is equal to 1.....

OpenStudy (anonymous):

(i) find k (ii) Evaluate P(X<6) ,P(x>=6) (iii) if P(X<=C) find the minimum value of C (iv)P[1.5<x<4.5 / x>2]

OpenStudy (amistre64):

i found k, its up there in the question box .....

OpenStudy (anonymous):

then 0 +k +2k +2k + 3k + k^2 +2k^2 +7k^2 + k=1

OpenStudy (anonymous):

by adding them then you can find k eassily....

OpenStudy (anonymous):

yeah!!

OpenStudy (amistre64):

.. why did i think that was going to infinity :/ good eye

OpenStudy (anonymous):

P(x < 6) = P(x=0) + P (x = 1) + P (x = 2) + P (x = 3) + P (x = 4) + P (x = 5) (Or) = 1 − P(x ≥ 6) = 1 − [P(x=6) + P(x=7)]

OpenStudy (anonymous):

if P(X<=C) is nothing but for C what value minimum we can take...